three men do a job in 12 days.Two of the men work twice as fast as third.How long will it take one of the faster men.
Answer by richwmiller(17219)

1/2x+1/x+1/x=1/12
1/2x+2/x=1/12
1/2x+4/2x=1/12
5/(2x)=1/12
5/x=2/12
x=30 faster
check
1/60+1/30+1/30=1/12
1/60+2/60+2/60=1/12
5/60=1/12
1/12=1/12
ok